OpenStudy (anonymous):

Divide p(x) by g(x),where p(x) = x + 3x^2 - 1 and g(x)=1+x

OpenStudy (anonymous):

3x-2 with remainder 1

OpenStudy (anonymous):

\[p(x)=x + 3x ^{2} - 1 and g(x) = 1+x\]

OpenStudy (anonymous):

its 3 x square

OpenStudy (anonymous):

i understand that ...!! i am getting remainder as 1.with q=3x-2

OpenStudy (anonymous):

but i didn't get u

OpenStudy (anonymous):

Quotient=3x-2 and remainder =1, still? :O

OpenStudy (anonymous):

ok i understood

OpenStudy (anonymous):

thankx

OpenStudy (anonymous):

Anytime :)